OpenStudy (anonymous):

Consider the function. y = 13 e^( x) At what point does the curve have maximum curvature?

OpenStudy (anonymous):

you want i guess i would say at 0

OpenStudy (anonymous):

\[\frac{\sqrt{(1+f')^3}}{f''}\] i think but i could be wrong

OpenStudy (anonymous):

well I used the formula |f"(x)|/[1+(f'(x)^2]^ 3/2 and got 13e^x/[1+(13e^x)^2]^3/2 then differentiated quotient rule [1+169^2x]^3/2.13e^x-13e^x.[3/2(1+169e^2x)^1/2.169e^2x]/[1+169e^2x]^3 now I have hard time setting the numerator to 0

OpenStudy (anonymous):

maybe i am wrong but i think your formula is upside down. let me check

OpenStudy (anonymous):

yeah your formula is upside down

OpenStudy (anonymous):

i cheated and took derivative here http://www.wolframalpha.com/input/?i=sqrt%28%281%2B%2813e^x%29^2%29%29^3%2F%2813e^x%29

OpenStudy (anonymous):

oh i see you want max and i am finding min, but min of radius of curvature is max of curvature, so it amount to the same thing. we set \[338e^{2x}-1=0\] solve for x to get it

OpenStudy (anonymous):

since all that other stuff in the derivative is never zero

OpenStudy (anonymous):

Thank you satellite73!

OpenStudy (anonymous):

A solution and plot from Mathematica is attached.

OpenStudy (anonymous):

Thank you as well robtobey!

OpenStudy (anonymous):

let me see if it is the same.. \[338e^{2x}=1\] \[e^{2x}=\frac{1}{338}\] \[2x=\ln(\frac{1}{338})=-\ln(338)\] \[x=-\frac{\ln(338)}{2}\]\]